OpenStudy (anonymous):

Solve the inequality and graph the solution set on a number line. Be sure to use closed dots or open circles. Do not use parentheses or brackets. (x+6)/(x−5) ≤ 0

OpenStudy (anonymous):

if u want me to answer ur quesion give me a medal first! KK

OpenStudy (anonymous):

u see how sweat i am i gave u a medal for ur question!

OpenStudy (anonymous):

@anisah I hope you're not serious. This is a disgusting practice. Not only are medals near-worthless in utility, this site was built on the principle of free help. @orbie Do out (x+6)/(x-5)=1+11/(x-5)≤0,11/(x-5)≤-1,(x-5)/11≥-1,x-5≥-11,x≥-6, which is a solid dot on x=-6 extending out in the positive direction.

OpenStudy (anonymous):

@badreferences talk abt ur self and urself only k dude

OpenStudy (anonymous):

@anisah I'm pretty sure it's tacitly understood that one should be a decent human being.

OpenStudy (anonymous):

@badreferences thank you for replying, I don't understand how you got the answer. Can you clarify please? =]

OpenStudy (anonymous):

I'm sorry, @orbie, I made a mistake. I did algebra wrong. XD Don't submit the answer yet, let me correct it.

OpenStudy (anonymous):

Sorry for taking a while to respond. XD Let me try typing that again, the HTML messed up on my screen.

OpenStudy (anonymous):

(x+6)/(x-5)≤0 means it's either zero or negative. It's only zero at x=-6, since then it'd be 0/(stuff)=0. When is it negative? It's negative when x+6>0 and x-5<0, or x+6<0 and x-5>0. In the former case, x>-6 and x<5; in the latter x<-6 and x>5. For no value is the latter true, so we can say x>-6 and x<5... BUT it can be evaluated at x=-6, but not at x=5, thus -6≤x<5, which is a solid point at x=-6, a hollow one at x=5, and a solid line in between.

OpenStudy (anonymous):

Thank you so much!